# DreamSteerer: Enhancing Source Image Conditioned Editability using Personalized Diffusion Models

### v1

@@ 1 Introduction
This challenge has sparked extensive research in T2I Personalization [16, 62, 10]. Specifically, personalization is the process of teaching T2I DPMs a novel visual concept with a few (usually 3-8) reference images by fine-tuning the pre-trained model parameters. After personalization, the personal concept is linked to a rare token in the text encoder dictionary, e.g., "[my_dog]", enabling flexible reuse of the visual concept in new contexts, e.g., "a photo of [my_dog] wearing an astronaut suit".
Despite the success of these models, the majority of previous works have been focusing on personalized generation, emphasizing the preservation of concept identity across varying textual conditions. In this work, we explore a natural extension on these methods to perform image editing with acquired concepts, namely personalized editing, which can be promising in enabling higher level control over the editing direction than traditional text-driven editing frameworks [55, 34, 21]. The primary goal is to synthesize a high-fidelity image that aligns the appearance and content of the target concept, as well as the structural layout and background of the source image without blindly resorting to copying learned reference images. A straightforward solution to this problem is to incorporate a personalized diffusion model into an existing text-driven editing model, by which personalized editing appears to be trivial via swapping a source subject token in the source prompt by the special token, e.g., "a photo of (dog $\rightarrow$ [my_dog]) wearing an astronaut suit". However, such a naïve incorporation often leads to severe distortion or failure in natural adaptation to the source image layout. We attribute the essential causes for such failure to the limited scope of reference images in personalization. Under a lack of data diversity, existing personalization methods are prone to collapsing into the patterns of reference images [71] and entangling subject-relevant and subject-irrelevant information [9]. This causes a significant loss in the model’s prior knowledge related to the source category, resulting in poor editability [60, 48] in a new context. Additionally, the stricter demands for maintaining structural layout [54, 72] and preserving subject-irrelevant information [21, 12] necessitate a higher level of editability than personalized generation. In more challenging editing scenarios, a certain level of extrapolation on the attributes of personalized concepts may be required, e.g. the editing process may require significant change in subject structure, pose or style to achieve a high-fidelity editing result on the source image. Without information on the source image during personalization, existing personalization methods can hardly be guaranteed to be editable on arbitrary source images.
@@ 3 Preliminaries
Existing text-driven editing works like DDS generate high-quality output for image editing with open-world vocabulary; however, their capability of personalized concepts editing remains underexplored. In this work, we focus on bridging such a gap. Given a personalized diffusion model $\epsilon_{\phi}$ trained via personalization in Eq. 2 and the latent state of a source image ${\mathbf{x}}^{\rm src}_{0}$ with an aligned source prompt ${\mathbf{y}}^{\mathrm{src}}$ (e.g., "a photo of a cat sitting next to a mirror"), our objective is to edit ${\mathbf{x}}^{\rm src}_{0}$ using the personalized concept $[s]$ captured by $\epsilon_{\phi}$. For example, an edit might be "a photo of a (cat $\rightarrow[s]$) sitting next to a mirror" where $[s]$ represents your childhood pet cat. We define the target prompt as ${\mathbf{y}}^{\rm ref}$. The desiderata for personalized editing are concluded as follows
@@
3.1 Preliminary Editing with Existing Personalized Baseline
@@ 4 The DreamSteerer Method
Figure 3: Overal framework of DreamSteerer (the gradient flows are illustrated with dashed lines).
@@ 4.1 Editability Driven Score Distillation (EDSD)
Then, we define the editability-driven loss as $\mathcal{L}_{\mathrm{ED}}=\mathbb{E}_{t,\epsilon}[\|\epsilon_{\phi}\left(\hat{x}_{t}(\phi),{\mathbf{y}}^{\mathrm{ref}},t\right)-\epsilon_{\phi_{0}}\left({\mathbf{x}}_{t}^{\rm src},{\mathbf{y}}^{\mathrm{src}},t\right)\|_{2}^{2}]$ which is similar to the $\mathcal{L}_{\rm DDS-S}$ but with different parameters to optimize. To reduce the computational cost, with a slight abuse of notation, we use $\nabla_{\phi}\mathcal{\mathcal{missing}}{L}_{\mathrm{ED}}\approx\frac{\partial\mathcal{L}_{\rm ED}}{\partial\hat{x}}\frac{\partial\hat{x}}{\partial\phi}$
@@ 7 Acknowledgement
This research was, in part, funded by the U.S. Government – DARPA ECOLE HR00112390061 and DARPA TIAMAT HR00112490421. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government.
@@ Appendix C More Implementation details
The pre-trained checkpoints provided by ViCo [20] has 16 concepts, which include 6 toys, 5 live animals, 2 types of accessories, 2 types of containers, and 1 building. The corresponding text prompts of source images are generated by BLIP2 [41]. For editing with one-shot personalization, performance evaluations are conducted on a subset of 20 cat and dog images.
@@
The EDSD gradient in Eq. 7 includes a Jacobian term $\frac{\partial\epsilon_{\phi}\left(\hat{x}_{t}(\phi),{\mathbf{y}}^{\mathrm{ref}}\right)}{\partial\hat{x}_{t}(\phi)}$, which is computationally expensive and unstable for optimization. Previous work [57] shows that this term can cause instability in optimization and omits the computation of it by setting it to ${\mathbf{I}}$. As shown in Fig. 8 we find that by setting it to ${\mathbf{I}}$ following previous work, the editing process tends to destroy the structural layout and background of the source image. Meanwhile the edited image tends to be over-saturated. This indicates that by setting $\frac{\partial\epsilon_{\phi}\left(\hat{x}_{t}(\phi),{\mathbf{y}}^{\mathrm{ref}}\right)}{\partial\hat{x}_{t}(\phi)}={\mathbf{I}}$, the discrepancy between personalized and source model score estimations has been maximized. We find that setting $\frac{\partial\epsilon_{\phi}\left(\hat{x}_{t}(\phi),{\mathbf{y}}^{\mathrm{ref}}\right)}{\partial\hat{x}_{t}(\phi)}=-{\mathbf{I}}$ leads to significantly better results with natural adaptation to the layout of the source image. We suspect that this is because, in the former work [57], the input noisy latent state is independent from the diffusion model. However, in Eq. 7, there exists an entanglement between negative score estimation in the single-step denoising function (Eq. 6) and the score estimation. We leave a further investigation on this interesting scenario to future study.
@@ Appendix E More ablation study
We ablate on two hyper-parameters $\lambda$ and cfg introduced in Eq. 9 for spatial feature guided sampling. As shown in Fig. 13, in contrast to usual selection of large cfg, e.g., 7 or 15. A relatively small cfg of 3.5 leads to better naturalness of the generated samples that avoids obvious subject appearance drift from the reference images. Meanwhile, a large spatial feature guidance scale $\lambda$ produces samples that better preserve the structure of source image. Thus we set cfg as 3.5 and $\lambda$ as 15.
@@ Appendix F More results compared with the base personalization methods
More results compare with the baseline models are provided in Fig. 14 15 16.
@@ Appendix I Explanation on editing direction visualization
As shown in Fig. 22, we observe that the attention maps from decoder layers with resolutions $16$ and $32$ have accurate subject layout, Thus we use the averaged CA maps from these layers as the subject mask. Similar to the spatial feature extraction process in Sec. 4.2, DDIM inversion is conducted on the source image ${\mathbf{x}}^{\rm src}_{0}$ by which the averaged CA map that corresponding to the source subject token is extracted, i.e., $M({\mathbf{x}}^{\rm src}_{0})=\mathbb{E}_{t,\ell}\operatorname{Softmax}\left(Q^{\ell}_{t}\left(K^{\ell}_{t}\right)^{T}/{\sqrt{C^{\ell}}}\right)$.
@@ 43. Answer: [Yes]
  44. Justification: We have thoroughly go through the NeurIPS Code of Ethics and confirm that the research conducted in this work conforms with it in every aspect.
@@ 48. Answer: [Yes]
  49. Justification: We have provide a discussion on potential social impact of the work in Sec. M.

### v2

@@ 1 Introduction
This challenge has sparked extensive research in T2I Personalization [16, 62, 10]. Specifically, personalization is the process of teaching T2I DPMs a novel visual concept with a few (usually 3-8) reference images by fine-tuning the pre-trained model parameters. After personalization, the personal concept is linked to a rare token in the text encoder dictionary, e.g., "[my_dog]", enabling flexible reuse of the visual concept in new contexts, e.g., "a photo of [my_dog] wearing an astronaut suit". Despite the success of these models, the majority of previous works have been focusing on personalized generation, emphasizing the preservation of concept identity across varying textual conditions. In this work, we explore a natural extension on these methods to perform image editing with acquired concepts, namely personalized editing, which can be promising in enabling higher level control over the editing direction than traditional text-driven editing frameworks [55, 34, 21]. The primary goal is to synthesize a high-fidelity image that aligns the appearance and content of the target concept, as well as the structural layout and background of the source image without blindly resorting to copying learned reference images. A straightforward solution to this problem is to incorporate a personalized diffusion model into an existing text-driven editing model, by which personalized editing appears to be trivial via swapping a source subject token in the source prompt by the special token, e.g., "a photo of (dog $\rightarrow$ [my_dog]) wearing an astronaut suit". However, such a naïve incorporation often leads to severe distortion or failure in natural adaptation to the source image layout. We attribute the essential causes for such failure to the limited scope of reference images in personalization. Under a lack of data diversity, existing personalization methods are prone to collapsing into the patterns of reference images [71] and entangling subject-relevant and subject-irrelevant information [9]. This causes a significant loss in the model’s prior knowledge related to the source category, resulting in poor editability [60, 48] in a new context. Additionally, the stricter demands for maintaining structural layout [54, 72] and preserving subject-irrelevant information [21, 12] necessitate a higher level of editability than personalized generation. In more challenging editing scenarios, a certain level of extrapolation on the attributes of personalized concepts may be required, e.g. the editing process may require significant change in subject structure, pose or style to achieve a high-fidelity editing result on the source image. Without information on the source image during personalization, existing personalization methods can hardly be guaranteed to be editable on arbitrary source images.
@@ 3 Preliminaries
Existing text-driven editing works like DDS generate high-quality output for image editing with open-world vocabulary; however, their capability of personalized concepts editing remains underexplored. In this work, we focus on bridging such a gap. Given a diffusion model $\epsilon_{\phi}$ personalized via Eq. 2 and the latent state of a source image ${\mathbf{x}}^{\rm src}_{0}$ with an aligned source prompt ${\mathbf{y}}^{\mathrm{src}}$ (e.g., "a photo of a cat sitting next to a mirror"), our objective is to edit ${\mathbf{x}}^{\rm src}_{0}$ using the personalized concept $[s]$ captured by $\epsilon_{\phi}$. For example, an edit might be "a photo of a (cat $\rightarrow[s]$) sitting next to a mirror," where $[s]$ represents your childhood pet cat. We define the target prompt as ${\mathbf{y}}^{\rm ref}$. The desiderata for personalized editing are concluded as follows
@@
3.1 Preliminary Editing with Existing Personalization Baseline
@@ 4 The DreamSteerer Method
Figure 3: Overall framework of DreamSteerer (the gradient flows are illustrated with dashed lines).
@@ 4.1 Editability Driven Score Distillation (EDSD)
Then, we define the editability-driven loss as $\mathcal{L}_{\mathrm{ED}}=\mathbb{E}_{t,\epsilon}[\|\epsilon_{\phi}\left(\hat{x}_{t}(\phi),{\mathbf{y}}^{\mathrm{ref}},t\right)-\epsilon_{\phi_{0}}\left({\mathbf{x}}_{t}^{\rm src},{\mathbf{y}}^{\mathrm{src}},t\right)\|_{2}^{2}]$, which is similar to the $\mathcal{L}_{\rm DDS-S}$ but with different parameters to optimize. To reduce the computational cost, with a slight abuse of notation, we use $\nabla_{\phi}\mathcal{\mathcal{missing}}{L}_{\mathrm{ED}}\approx\frac{\partial\mathcal{L}_{\rm ED}}{\partial\hat{x}}\frac{\partial\hat{x}}{\partial\phi}$
@@ 7 Acknowledgement
Special thanks to Peter Tu for providing us with his valuable inputs and support that greatly helped improving our work and for motivating and encourage us on further implementations of our work. This research was, in part, funded by the U.S. Government – DARPA ECOLE HR00112390061 and DARPA TIAMAT HR00112490421. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government.
@@ Appendix C More Implementation details
The pre-trained checkpoints provided by ViCo [20] have 16 concepts, which include 6 toys, 5 live animals, 2 types of accessories, 2 types of containers, and 1 building. The corresponding text prompts of source images are generated by BLIP2 [41]. For editing with one-shot personalization, performance evaluations are conducted on a subset of 20 cat and dog images.
@@
The EDSD gradient in Eq. 7 includes a Jacobian term $\frac{\partial\epsilon_{\phi}\left(\hat{x}_{t}(\phi),{\mathbf{y}}^{\mathrm{ref}}\right)}{\partial\hat{x}_{t}(\phi)}$, which is computationally expensive and unstable for optimization. Previous work [57] shows that this term can cause instability in optimization and omits the computation of it by setting it to ${\mathbf{I}}$. As shown in Fig. 8 we find that by setting it to ${\mathbf{I}}$ following previous work, the editing process tends to destroy the structural layout and background of the source image. Meanwhile, the edited image tends to be over-saturated. This indicates that by setting $\frac{\partial\epsilon_{\phi}\left(\hat{x}_{t}(\phi),{\mathbf{y}}^{\mathrm{ref}}\right)}{\partial\hat{x}_{t}(\phi)}={\mathbf{I}}$, the discrepancy between personalized and source model score estimations has been maximized. We find that setting $\frac{\partial\epsilon_{\phi}\left(\hat{x}_{t}(\phi),{\mathbf{y}}^{\mathrm{ref}}\right)}{\partial\hat{x}_{t}(\phi)}=-{\mathbf{I}}$ leads to significantly better results with natural adaptation to the layout of the source image. We suspect that this is because, in the former work [57], the input noisy latent state is independent from the diffusion model. However, in Eq. 7, there exists an entanglement between negative score estimation in the single-step denoising function (Eq. 6) and the score estimation. We leave a further investigation on this interesting scenario to future study.
@@ Appendix E More ablation study
We ablate on two hyper-parameters $\lambda$ and cfg introduced in Eq. 9 for spatial feature guided sampling. As shown in Fig. 13, in contrast to the usual selection of large cfg, e.g., 7 or 15. A relatively small cfg of 3.5 leads to better naturalness of the generated samples that avoids obvious subject appearance drift from the reference images. Meanwhile, a large spatial feature guidance scale $\lambda$ produces samples that better preserve the structure of the source image. Thus we set cfg as 3.5 and $\lambda$ as 15.
@@ Appendix F More results compared with the base personalization methods
More results compared with the baseline models are provided in Fig. 14 15 16.
@@ Appendix I Explanation on editing direction visualization
As shown in Fig. 22, we observe that the attention maps from decoder layers with resolutions $16$ and $32$ have accurate subject layout. Thus, we use the averaged CA maps from these layers as the subject mask. Similar to the spatial feature extraction process in Sec. 4.2, DDIM inversion is conducted on the source image ${\mathbf{x}}^{\rm src}_{0}$, by which the averaged CA map that corresponds to the source subject token is extracted, i.e., $M({\mathbf{x}}^{\rm src}_{0})=\mathbb{E}_{t,\ell}\operatorname{Softmax}\left(Q^{\ell}_{t}\left(K^{\ell}_{t}\right)^{T}/{\sqrt{C^{\ell}}}\right)$.
@@ 43. Answer: [Yes]
  44. Justification: We have thoroughly gone through the NeurIPS Code of Ethics and confirm that the research conducted in this work conforms with it in every aspect.
@@ 48. Answer: [Yes]
  49. Justification: We have provided a discussion on the potential social impact of the work in Sec. M.